A person climbs to the top of a mountain at the rate of 1.5 kilometers per hour and down the same way at a rate of 4.5 kilometers per hour.If he is gone for 6 hours, how far did he travel from camp to the top of the mountain?
Answer by checkley77(12569)

D=RT
D=1.5T UP THE MOUNTAIN.
D=4.5(6-T) DOWN THE MOUNTAIN
SET THE DISTANCES EQUAL & SOLVE FOR THE TIME.
1.5T=4.5(6-T)
1.5T=27-4.5T
1.5T+4.5T=27
6T=27
T=27/6
T=4.5 HOURS FOR THE TIME GOING UP THE MOUNTAIN.
D=1.5*4.5=6.75 MILES IS THE DISTANCE UP THE MOUNTAIN.
PROOF:
6.75=4.5(6-4.5)
6.75=4.5*1.5
6.75=6.75